# Beyond Human Norms: Unveiling Unique Values of Large Language Models through Interdisciplinary Approaches

### v1

@@ 3.1 Formalization and Overview
In this work, we aim to establish the fundamental shared value dimensions of a wide range of LLMs and then evaluate their inclinations towards these values. Define $p_{\theta}(y|x)$ as an LLM parameterized by $\theta$ which generates a response $y\in\mathcal{Y}$ from a given input question $x\in\mathcal{X}$, where $\mathcal{Y}$ and $\mathcal{X}$ are the spaces of all responses and questions. By considering each LLM as an AI respondent (participant), we incorporate a set of diverse respondents $\mathcal{P}=\{p_{{}_{1}},\dots\,p_{N}\}$ across various model sizes, training methods, and data. The core components of ValueLex are: Generative Value Construction and Projective Value Evaluation.
@@ 3.2 ValueLex: Generative Value Construction
Step 1: Value Elicitation  Besides including various LLMs with different architectures and training methods, such as LLaMA (Touvron et al., 2023b), GPT (Ouyang et al., 2022), Mistral (Jiang et al., 2023) and ChatGLM (Zeng et al., 2022), we also assign different configurations $\theta_{k}\in\Theta$ to each LLM $p_{i}$, like decoding temperature and probability threshold of top-$p$ sampling, where $\Theta$ is the space of all valid configurations, to increase respondent diversity. At last, we get $N=525$ respondents in total. More details of LLM respondents are given in Appendix 2. Since there are no well-designed psychometrics inventories and to mitigate the noise inherent in LLM responses (Westland, 2022; Li et al., 2023b), we eschew traditional Likert scaling questionnaires. Instead, we elicit values by asking LLMs to respond to carefully designed and diverse questions $x\in\mathcal{Q}$ with similar meanings, e.g., $q=$If my responses are based on certain values, the terms are as follows”, and obtain a set of candidate value descriptors $\mathcal{V}=\{v_{1},v_{2},\ldots\}$, which is formulated as:
@@
in which we leverage word embeddings to measure similarity. Once getting $K$ clusters, we employ a trained LLM to semantically induce the most fitting name $v$ for each cluster according to the candidates in it, to eschew subjective biases, and obtain the final value descriptors $V=\{v_{1},v_{2},\ldots,v_{K}\}$ that reflect LLMs’ ethical and moral compass.
@@ 3.3 ValueLex: Value Evaluation via Projective Test
For $\phi$, we adapt a similar six-scale scoring system as in (Rotter, 1950), that is, 6 indicates a positive alignment with the value dimension, 3 is neutral, and 0 signifies a conflict. In practice, we manually score a small set of responses and instantiate $\phi$ with GPT-4, which demonstrated similar performance as human annotators (Gilardi et al., 2023), in a few shot chain-of-though manner (Wei et al., 2022b), and then normalize the scores into $[0,1]$. In our experiments, the Quadratic Weighted Kappa between human and $\phi$ is 0.8, indicating the reliability of our implementation. The whole process is outlined in Algorithm 2.

### v2

@@ 3.1 Formalization and Overview
In this work, we aim to establish the fundamental shared value dimensions of a wide range of LLMs and then evaluate their inclinations towards these values. Define $p(y|x)$ as an LLM which generates a response $y\in\mathcal{Y}$ from a given input question $x\in\mathcal{X}$, where $\mathcal{Y}$ and $\mathcal{X}$ are the spaces of all responses and questions. By considering each LLM as an AI respondent (participant), we incorporate a set of diverse respondents $\mathcal{P}=\{p_{{}_{1}},\dots\,p_{N}\}$ across various model sizes, training methods, and data. The core components of ValueLex are: Generative Value Construction and Projective Value Evaluation.
@@ 3.2 ValueLex: Generative Value Construction
Step 1: Value Elicitation  Besides including various LLMs with different architectures and training methods, such as LLaMA (Touvron et al., 2023b), GPT (Ouyang et al., 2022), Mistral (Jiang et al., 2023) and ChatGLM (Zeng et al., 2022), we also assign different configurations $\theta_{k}\in\Theta$ to each LLM $p_{i}$, like decoding temperature and probability threshold of top-$p$ sampling, where $\Theta$ is the space of all valid configurations, to increase respondent diversity. At last, we get $N=525$ respondents in total. More details of LLM respondents are given in Appendix 2. Since there are no well-designed psychometrics inventories and to mitigate the noise inherent in LLM responses (Westland, 2022; Li et al., 2023b), we eschew traditional Likert scaling questionnaires. Instead, we elicit values by asking LLMs to respond to carefully designed and diverse questions $q\in\mathcal{Q}$ with similar meanings, e.g., $q=$If my responses are based on certain values, the terms are as follows”, and obtain a set of candidate value descriptors $\mathcal{V}=\{v_{1},v_{2},\ldots\}$, which is formulated as:
@@
in which we leverage word embeddings to measure similarity. Once getting $K$ clusters, we employ a trained LLM to semantically induce the most fitting name for each cluster according to the candidates in it, to eschew subjective biases, and obtain the final value descriptors $V=\{v_{1},v_{2},\ldots,v_{K}\}$ that reflect LLMs’ ethical and moral compass.
@@ 3.3 ValueLex: Value Evaluation via Projective Test
For $\phi$, we adopt a similar six-scale scoring system as in (Rotter, 1950), that is, 6 indicates a positive alignment with the value dimension, 3 is neutral, and 0 signifies a conflict. In practice, we manually score a small set of responses and instantiate $\phi$ with GPT-4, which demonstrated similar performance as human annotators (Gilardi et al., 2023), in a few shot chain-of-though manner (Wei et al., 2022b), and then normalize the scores into $[0,1]$. In our experiments, the Quadratic Weighted Kappa between human and $\phi$ is 0.8, indicating the reliability of our implementation. The whole process is outlined in Algorithm 2.
@@ A.3 Value Descriptors of PLMs and Aligned LLMs
The following words represent a subset of the 197 descriptors elicited from aligned models and 564 from pretrained ones.
Aligned Descriptors Pretrained Descriptors
Accuracy Religion
Friendliness Family
Empathy Leisure
Inclusivity Freedom
Integrity Faith
Justice Independence
Respect Power
Transparency Cherish
Trustworthiness Environment
Honesty Expression
Creativity Innovation
Collaboration Sustainability
Efficiency Empowerment
Professionalism Visionary
Kindness Community
Altruism Money
Fairness Leadership
Unbiased Uniqueness
Factual Art
User-friendly Law
Helpful Marriage
Equality Nonconformity
Clear Work
Accessible Loyalty